Accelerator Update for February 27, 2009 - March 2, 2009 |
The Integrated Luminosity for 2/23/09 to 3/2/09 was 49.17 Pb-1. Friday Feb 27 The day shift began with the Tevatron (TeV), Recycler, Linac, Booster, and the Main Injector (MI) recovering from the Kautz Road power glitch, with Operations monitoring the Antiproton Source stack (24.1E10), and with the MiniBooNE, NuMI, the Collider Detector at Fermilab (CDF), and D-Zero (D0) experiments waiting for beam. Operations began sending beam to MiniBooNE around 8 AM. Operations reported at 8:47 AM that they had successfully hipotted the TeV. A controls expert reported at 8:51 AM that they had installed a PLC to control the I- Source. A water expert reported at 8:55 AM that work had begun on the MI-40 LCW pump #2. This work took about an hour to complete. Operations reported at 9:38 AM that operators had rebooted the MI vacuum front end. Operations prepared D0 for a one-hour access at 10:04 AM. At 10:59 AM, Operations reported that the Booster RF system was sparking so much that operators had to lower Booster’s beam intensity. Linac experts reported at 11:13 AM that they were working on LRF5. The Neutron Therapy Facility reported at 11:25 AM that that they had completed treatment on their last patient for the day. At 11:44 AM, a RF expert reported a bad pulsed power breaker circuit at MI-60 that provided power to the bias power supplies. They will try using the test stand breaker temporarily. Operations prepared D0 for another one-hour access at 12:18 PM. Operations put the TeV through a dry squeeze at 2:29 PM. Operations began sending beam to NuMI around 2:50 PM. At 3:59 PM, Operations reported having problems getting beam to NuMI due to beam in the MI kicker gap. Experts investigated. Operations put the TeV through a wet squeeze at 4:34 PM. At 6:28 PM, Operators accessed the Pbar tunnels and found a LCW leak from a magnet (D6Q3). Operators isolated the magnet and called in experts to repair the leak. At 6:30 PM, a RF expert reported at problem with the Accumulator 4-8 GHz core momentum cooling system. He was investigating the light link system. Pbar resumed stacking at 8:43 PM. At 9:02 PM, Operations halted stacking so that experts could conduct further work on the Accumulator 4-8 GHz momentum system. Pbar resumed stacking at 10:04 PM. Operations reported at 11:59 PM that the Recycler had stashed antiprotons once during the evening shift. Saturday Feb 28 The midnight shift began with TeV experts conducting studies, with Operations monitoring the stash (69.0E10), with Pbar stacking (21.9E10), and with NuMI and MiniBooNE taking beam. TeV studies ended at 2:59 AM. Operations put the TeV through a dry squeeze at 3:07 AM. Operations put the TeV into shot setup at 3:59 AM. Operations established store 6840 at 5:11 AM with an initial luminosity of 157.2E30. At 6:15 AM, Operations reported that the first transfer of antiprotons was lost during a stash. The rest of the transfers went okay. At 7:04 AM, LRF3 tripped off. It reset. Operations reported at 8:05 AM that the Recycler had stashed eight times during the midnight shift. At 9:05 AM, the P1 beam permit tripped off when a power supply’s current (I:V714) drifted low. Everything reset. Operations terminated store 6840 at 4:37 PM. Operations prepared CDF for an access at 4:40 PM. CDF came out of access at 7:46 PM. Operations reported at 8 PM that the Recycler had stashed 12 times during the last 12 hours. Operations put the TeV through a dry squeeze at 8:05 PM. Operations put the TeV through a wet squeeze at 8:27 PM. Operations put the TeV into shot setup at 9:42 PM. At 10:59 PM, the P1 transfer line tripped off when I:V714’s current drifted low. Everything reset. Operations established store 6844 at 11:02 PM with an initial luminosity of 316.91E30. Sunday March 1 The midnight shift began with Operations monitoring store 6844 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. At 4:05 AM, LRF3 tripped off. It reset. At 5:17 AM, I:V714 caused the P1 transfer line to trip off. Everything reset. Operations reported at 8:01 AM that the Recycler had stashed 11 times during the last 12 hours. Operations terminated store 6844 at 9:20 AM. Operations prepared CDF for a two-hour access at 9:30 AM. Operations put the TeV through a wet squeeze at 12:21 PM. At 1:09 PM, Operations reported a kicker prefire (PAK5), which aborted the protons. Operations put the TeV through a dry squeeze at 1:12 PM. Operations put the TeV through wet squeeze at 1:59 PM. At 2:24 PM, LRF3 tripped off. It reset. Operations put the TeV through a wet squeeze at 2:45 PM. Operations put the TeV into shot setup at 3:23 PM. At 3:51 PM, Operations reported that the MI Power Supply interlock Loop had tripped off three times and then tripped off again about 15 minutes later. Operations contacted an expert. Operations reported at 4:55 PM that MI was up and ramping. Operations established store 6850 at 5:59 PM with an initial luminosity of 319.98E30. Operations reported at 8 PM that the Recycler had stashed eight times during the last 12 hours. At 8:25 PM, BRF4 tripped off for the first of two times in 14 minutes. It reset. At 9:58 PM, BRF15 tripped off. It reset. At 9:59 PM, BRF3 and 4 tripped off for the first of two times in 15 minutes. They reset. At 10:14 PM, BRF3, 4, and 9 tripped off for the first of three times in two minutes. They reset. At 10:22 PM, BRF15 and 16 tripped off. They reset. At 10:24 PM, BRF3 and 8 tripped off. They reset. At 10:25 PM, Operations reported that due to the Booster RF sparking trips, they were lowering the Booster beam intensity. Operations tuned the accelerator and then brought it back up to normal intensity about 20 minutes later. At 10:59 PM, Operations reported a problem with a Pbar Debuncher quadrupole power supply (D: QD). At 11:16 PM, Operations reported that D:QD had no control power. An operator called in an expert. At 11:57 PM, BRF3 tripped off for the first of three times in about 30 minutes. It reset. Monday March 2 The midnight shift began with Operations monitoring store 6850 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. At 12:24 AM, BRF15 tripped off. It reset. At 3:04 AM, BRF4 tripped off for the first of two times in about 50 minutes. It reset. At 4:08 AM, BRF8 tripped off. It reset. At 4:35 AM, BRF9 tripped off for the first of two times in 11 minutes. It reset. At 7:21 AM, BRF4 and 9 tripped off. They reset. Operations reported at 7:59 AM that the Recycler had stashed seven times during the last 12 hours. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow TeV experts to conduct an end-of-store study. |
|
More Information Comments and Suggestions |
| last modified 03/02/2009 email Fermilab |
| Security, Privacy, Legal | |